# Image-Text-Image Knowledge Transferring for Lifelong Person Re-Identification with Hybrid Clothing States

### v1

@@ 4.1. Datasets and Evaluation Protocol
Table 1. The statistical information of our LReID-Hybrid training benchmarks.
@@
Market-1501 (Zheng et al., 2015) 1,501 32,668 6 $\checkmark$
MSMT17 (Wei et al., 2018) 4,101 126,441 15 $\checkmark$
LTCC (Qian et al., 2020) 152 17,119 12
[…]
2613
PRCC (Yang et al., 2019) 221 33,698 3
[…]
2613
Table 2. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 1). The results are reported after the last training phase. The best results are shown in bold.
@@
We also evaluate our proposed method on the widely-used standard benchmarks involving only datasets without clothing changes (Wu and Gong, 2021; Ge et al., 2022), where Market-1501 (Zheng et al., 2015), DukeMTMC-reID (Zheng et al., 2017), CUHK-SYSU (Xiao et al., 2017) and MSMT17 (Wei et al., 2018) are used sequentially for lifelong learning. Note that DukeMTMC-reID is only used for academic use and fair comparison without identifying or showing pedestrian images. CUHK-SYSU is modified from the original for person search and rearranged following (Wu and Gong, 2021). For other datasets, we follow the training and evaluation protocols proposed in their original papers.
Unseen datasets. To evaluate the generalization of a model in the same-cloth scenario, we select five conventional ReID datasets, including CUHK01 (Li et al., 2013), CUHK02 (Li and Wang, 2013), GRID (Loy et al., 2010), SenseReID (Zhao et al., 2017) and PRID (Hirzer et al., 2011). We follow their standard protocols to conduct the query and gallery sets for testing. As for the cloth-changing scenario, we evaluate lifelong ReID models on VC-Clothes (Wan et al., 2020) and Celeb-reID (Huang et al., 2019), neither of which appeared in the training set. For the Celeb-reID dataset, we use its lightweight version (i.e., Celeb-reID-light) without losing generality, and each pedestrian in Celeb-reID-light does not wear the same clothes twice.
Evaluation metrics. We adopt mean Average Precision (mAP) and Rank-1 accuracy (R-1) for performance evaluation on seen domains so far after each training step. We also report the average accuracies by averaging mAP and R-1 in seen domains as well as the accuracies in unseen domains. Considering the great performance discrepancy between same-cloth and cloth-changing datasets and to evaluate the model’s ability in different clothing states, we report the average accuracies on seen same-cloth and cloth-changing datasets, respectively, which are denoted as $(\overline{s}_{\rm mAP}^{s},\overline{s}_{\rm R-1}^{s})$ and $(\overline{s}_{\rm mAP}^{c},\overline{s}_{\rm R-1}^{c})$. We denote the average accuracy on all unseen same-cloth and cloth-changing datasets as $(\overline{s}_{\rm mAP}^{us},\overline{s}_{\rm R-1}^{us})$ and $(\overline{s}_{\rm mAP}^{uc},\overline{s}_{\rm R-1}^{uc})$, respectively. For the LTCC, PRCC and VC-Clothes datasets, we evaluate in the cloth-changing setting where only cloth-changing samples are involved in the testing set.
4.2. Implementation Details
We adopt the CLIP (Radford et al., 2021) image encoder as the ReID model, which is ViT-B/16 (Dosovitskiy et al., 2021) with 12 Transformer (Vaswani et al., 2017) layers. The class token of the last layer is used as the visual representation of the input person image. The CLIP text encoder consists of 12 Transformer (Vaswani et al., 2017) layers and the [EOS] token of the last layer is considered as the text embedding given structured $P$ as input. The images are resized to 256 $\times$ 128 and augmented by random horizontal flipping, padding, cropping and erasing (Zhong et al., 2020). The batch size is set to 64, with 4 samples per pedestrian. For each domain, the model is optimized with Adam optimizer (Kingma and Ba, 2015) with weight decay of $1\times 10^{-4}$ in two stages. The first stage takes 120 epochs using a learning rate initialized as $3.5\times 10^{-4}$ with cosine learning rate decay. The second stage takes 60 epochs with the warmup strategy that linearly increases the learning rate from $5\times 10^{-7}$ to $5\times 10^{-6}$ in the first 10 epochs. The learning rate is then decreased by a factor of 10 at the 40th epoch. To achieve knowledge adaptation, we implement the slower-paced learner by decreasing the learning rate by a factor of 10 after the first domain. $M$ is set to 16. Following (Li et al., 2023), $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are set to 1, 0.25 and 1, respectively.
4.3. Comparison with the State-of-the-Art
Competitors. We first compare with a widely-used rehearsal-free method AKA (Pu et al., 2021) for LReID. Then, we reimplement all competitors using CLIP (Radford et al., 2021) pre-trained ViT-B/16 (Dosovitskiy et al., 2021) as the ReID model. Specifically, SFT denotes sequentially fine-tuning the model on each dataset, and Joint-Train denotes training on all datasets jointly. We also compare our proposed $Teata$ with LwF (Li and Hoiem, 2017) and CLIP-ReID (Li et al., 2023). LwF performs knowledge distillation loss on the logit values of old tasks to keep task-specific decision boundaries. CLIP-ReID leverages the CLIP (Radford et al., 2021) pre-trained text encoder and uses the learned text prompts to regularize the image encoder. CLIP-ReID is designed for conventional person ReID in a stationary domain, and we reimplement it to achieve lifelong learning by retraining text prompts on each sequentially coming dataset.
4.3.1. Results in Seen Domains
Without loss of generalization, we show the results of training with Order 1 and evaluating on seen datasets in Tab. 2. More results in other orders are provided in the appendix. We also show results on the standard LReID benchmarks in Tab. 3.
@@ 4.3.1. Results in Seen Domains
Table 3. Comparison with the state-of-the-art methods in the conventional LReID setting.w/ Ex.” denotes rehearsal-based methods using exemplars. The training order is Market-1501 $\rightarrow$ DukeMTMC-reID $\rightarrow$ CUHK-SYSU $\rightarrow$ MSMT17, and the results are reported after the last training phase. The best results are shown in bold.
@@
AKA (Pu et al., 2021) 59.7 80.1 32.7 48.3 82.0 84.4 17.1 34.9 47.9 61.9
GwFReID (Wu and Gong, 2021) $\checkmark$ 60.9 81.6 46.7 66.5 81.4 83.9 25.9 52.4 53.7 71.1
KRC (Yu et al., 2023) $\checkmark$ 60.3 82.3 58.7 72.7 88.9 90.5 43.3 67.7 62.8 78.3
PTKP (Ge et al., 2022) $\checkmark$ 75.8 89.7 62.0 76.7 85.0 86.3 34.5 60.9 64.3 78.4
LwF (Li and Hoiem, 2017) 54.0 76.2 64.7 80.7 85.6 87.5 57.7 80.1 65.5 81.1
CLIP-ReID (Li et al., 2023) 50.6 74.8 60.0 76.8 84.5 86.4 64.7 84.9 65.0 80.7
@@
Figure 3. Comparison results of our proposed method with SFT on each seen domain along the lifelong learning process. We report the results in Order 1 w.r.t. (a) mAP and (b) R-1.
@@
Results on LReID-Hybrid benchmarks. AKA (Pu et al., 2021) achieves relatively poor results in our proposed LReID-Hybrid task. Although AKA uses a knowledge graph to represent and accumulate knowledge, it still operates in the image space, making it hard to maintain and accumulate knowledge when different tasks exhibit significant differences, especially when clothing status changes. Due to the absence of data from previous domains, SFT suffers catastrophic forgetting greatly. LwF performs knowledge distillation between current and frozen old models to overcome catastrophic forgetting at the cost of sacrificing performance in the current domain. Due to the great discrepancy of acquired knowledge across domains caused by variations of clothing state, LwF cannot make a good balance between knowledge acquired from domains with different clothing states. Although leveraging text, CLIP-ReID (Li et al., 2023) has no obvious advantage compared with LwF which fine-tunes the pre-trained CLIP image encoder without the supervision of text. The results show that the potential of text is still not fully unleashed. On the contrary, our proposed method successfully takes advantage of text semantics to balance and accumulate knowledge from various domains and shows a significant superiority over them.
@@
It is also worth noting the performance of Joint-Train is traditionally regarded as the proximate performance upper bound of the lifelong person ReID model. Similarly, when we train the standard and cloth-changing datasets jointly, we find the model can preserve its capability in the same-cloth scenario. However, the results in the cloth-changing domains are far from satisfactory. We argue that the acquired knowledge shows great discrepancy due to various clothing states across domains. The data imbalance in distribution, as well as the amount between standard and cloth-changing datasets, would make the model unexpectedly biased toward the former when they are used jointly for training. The results further reveal the challenges faced in our proposed LReID-Hybrid task that considers clothing variations during the real-world application of person ReID models. In contrast to the image model, the text encoder is trained by a mass of data, building a compact and well-represented space. Our proposed Teata leverages the text space as an intermediary for knowledge transfer and accumulation, which can effectively alleviate the representation bias caused by data imbalance.
@@
In Fig. 3, we show the results in each seen domain along the lifelong learning process. Without the help of text semantics, SFT achieves inferior results in each domain. What’s worse, when adapting to new domains, it forgets previously acquired knowledge dramatically. Thanks to the consistent and compact text semantics, our proposed method can accumulate new knowledge while the loss of previously acquired knowledge is significantly reduced. We also observe a slight performance improvement on the LTCC dataset when training on PRCC at step 4 compared with step 3, and for our proposed method, the final R-1 at step 4 is even higher than that at step 2. We argue that since both PRCC and LTCC are cloth-changing datasets, the model tends to learn cloth-irrelevant identity representations on them. Therefore, the learned cloth-irrelevant identity clues on PRCC are useful in improving the discriminative ability of pedestrians with clothing changes on LTCC. Our proposed $Teata$ exhibits a smoother forgetting trend in performance, showing its effectiveness in transferring and accumulating knowledge.
@@
Table 4. Unseen domain comparisons with state-of-the-art models trained in the LReID-Hybrid setting (Order 1). The results are given by the trained model after the last training step. $\dagger$’ means the light version of Celeb-reID. The best results are shown in bold.
@@
Table 5. Unseen domain comparisons with state-of-the-art models trained in the conventional LReID setting. The training order is Market-1501 $\rightarrow$ DukeMTMC-reID $\rightarrow$ CUHK-SYSU $\rightarrow$ MSMT17. The results are given by the trained model after the last training step. $\dagger$’ indicates the light version of Celeb-reID. The best results are shown in bold.
@@
PTKP (Ge et al., 2022) 71.1 69.4 66.4 66.1 33.9 25.3 53.9 46.1 38.4 27.3 24.2 21.8 5.3 9.2 52.7 46.8 14.8 15.5
KRC (Yu et al., 2023) 79.1 78.0 70.0 69.0 33.6 25.5 51.3 44.0 51.8 40.6 20.2 17.5 6.3 12.2 57.2 51.4 13.3 14.9
LwF (Li and Hoiem, 2017) 76.5 74.7 67.2 66.7 36.8 28.0 56.2 47.0 58.0 46.4 35.0 33.5 6.6 13.2 58.9 52.6 20.8 23.4
CLIP-ReID (Yu et al., 2023) 77.2 75.6 65.3 64.9 35.4 25.3 56.7 48.5 53.3 40.5 36.4 34.9 6.0 11.4 57.6 51.0 21.2 23.2
@@
Results on cloth-consistent LReID benchmarks. Since our proposed method is capable of handling the challenging LReID-Hybrid task, it is expected to also work well when pedestrians do not change their clothes. As shown in Tab. 3, we also compare our proposed method with state-of-the-art lifelong person ReID methods on the standard LReID benchmarks. Exemplars are considered to be of vital importance in relieving the catastrophic forgetting problem, and replay-based methods (Wu and Gong, 2021; Yu et al., 2023; Ge et al., 2022) achieve better results than AKA (Pu et al., 2021) indeed. However, saving person images can cause privacy issues, which is not realistic in real-world applications, and can bring great resource consumption with the increase of training domains. We also implement LwF (Li and Hoiem, 2017) and CLIP-ReID (Li et al., 2023) on the cloth-consistent LReID benchmarks. With the superior capability of CLIP, they achieve better results than state-of-the-art replay-based methods (Wu and Gong, 2021; Yu et al., 2023; Ge et al., 2022) on the current domain, but the performance on previous domains is still far from satisfactory, showing the catastrophic forgetting problem is still not well dealt with. In contrast, our proposed method achieves a good balance across domains. With the help of consistent text semantics, our proposed method beats all competitors by a large margin without relying on impractical and inflexible exemplars.
4.3.2. Results in Unseen Domains

### v2

@@ IV-A1 Seen Datasets
We also evaluate our proposed method on the widely-used standard benchmarks involving only datasets without clothing changes [35, 2], where Market-1501 [42], DukeMTMC-reID [44], CUHK-SYSU [45] and MSMT17 [43] are used sequentially for lifelong learning. DukeMTMC-reID is only used for academic use and fair comparison without identifying or showing pedestrian images. CUHK-SYSU is modified from the original for person search and rearranged following [35]. For other datasets, we follow the training and evaluation protocols proposed in their original papers.
IV-A2 Unseen Datasets
To evaluate the generalization of a model in the same-cloth scenario, we select five conventional ReID datasets, including CUHK01 [46], CUHK02 [47], GRID [48], SenseReID [49] and PRID [50]. We follow their standard protocols to conduct the query and gallery sets for testing. As for the cloth-changing scenario, we evaluate lifelong ReID models on VC-Clothes [51] and Celeb-reID [52], neither of which appeared in the training set. For the Celeb-reID dataset, we use its lightweight version (i.e., Celeb-reID-light) without losing generality, and each pedestrian in Celeb-reID-light does not wear the same clothes twice.
TABLE I: The statistics of our LReID-Hybrid training benchmarks.
@@ IV-A2 Unseen Datasets
Market-1501 [42] 1,501 32,668 6 $\checkmark$
MSMT17 [43] 4,101 126,441 15 $\checkmark$
LTCC [6] 152 17,119 12
PRCC [8] 221 33,698 3
IV-A3 Evaluation Metrics
We adopt mean Average Precision (mAP) and Rank-1 accuracy (R-1) for performance evaluation on seen domains so far after each training step. We also report the average accuracies by averaging mAP and R-1 in seen domains as well as the accuracies in unseen domains. Considering the great performance discrepancy between same-cloth and cloth-changing datasets and to evaluate the model’s ability in different clothing states, we report the average accuracies on seen same-cloth and cloth-changing datasets, respectively, which are denoted as $(\overline{s}_{\rm mAP}^{s},\overline{s}_{\rm R-1}^{s})$ and $(\overline{s}_{\rm mAP}^{c},\overline{s}_{\rm R-1}^{c})$. We denote the average accuracy on all unseen same-cloth and cloth-changing datasets as $(\overline{s}_{\rm mAP}^{us},\overline{s}_{\rm R-1}^{us})$ and $(\overline{s}_{\rm mAP}^{uc},\overline{s}_{\rm R-1}^{uc})$, respectively. For the LTCC, PRCC, and VC-Clothes datasets, we evaluate in the cloth-changing setting where only cloth-changing samples are involved in the testing set.
IV-B Implementation Details
We adopt the CLIP [11] image encoder as the ReID model, which is ViT-B/16 [53] with 12 Transformer [54] layers. The class token of the last layer is used as the visual representation of the input person image. The CLIP text encoder consists of 12 Transformer [54] layers and the [EOS] token of the last layer is considered as the text embedding given structured prompts as input. The images are resized to 256 $\times$ 128 and augmented by random horizontal flipping, padding, cropping, and erasing [55]. The batch size is set to 64, with 4 samples per pedestrian. For each domain, the model is optimized with Adam optimizer [56] with weight decay of $1\times 10^{-4}$ in two stages. The first stage takes 120 epochs using a learning rate initialized as $3.5\times 10^{-4}$ with cosine learning rate decay. The second stage takes 60 epochs with the warmup strategy that linearly increases the learning rate from $5\times 10^{-7}$ to $5\times 10^{-6}$ in the first 10 epochs. The learning rate is then decreased by a factor of 10 at the 40th epoch. In our proposed KAP strategy, the slower-paced learner is adopted to reduce the learning rate by 10 times after the first domain. $M$ is set to 16. Following [12], $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 1, 0.25 and 1.
TABLE II: Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 1). The results are reported after the last training phase. The best results are shown in bold.
@@ IV-B Implementation Details
Figure 3: Comparison results of our proposed method with SFT on each seen domain along the lifelong learning process. We report the results in Order 1 w.r.t. (a) mAP and (b) R-1.
@@
TABLE III: Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 2). The results are reported after the last training phase. The best results are shown in bold.
Training Order 2 Market-1501 LTCC PRCC MSMT17 SC Average CC Average
AKA [1] 52.2 75.0 7.6 17.6 26.5 25.8 21.8 41.8 37.0 58.4 17.1 21.7
SFT 49.5 74.5 11.6 25.3 22.3 20.8 63.8 84.7 56.7 79.6 22.8 23.1
LwF [33] 58.7 79.6 14.2 27.6 32.1 32.0 60.2 82.1 59.5 80.9 23.2 29.8
CLIP-ReID [12] 56.3 80.0 11.9 26.0 25.2 26.2 67.6 86.5 62.0 83.3 18.9 26.1
Teata (Ours) 82.9 93.0 17.7 35.2 47.4 45.9 65.9 85.0 74.4 89.0 32.6 40.6
Joint-Train 89.1 95.0 13.6 28.8 45.8 47.6 70.3 86.6 79.7 90.8 29.7 38.2
TABLE IV: Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 3). The results are reported after the last training phase. The best results are shown in bold.
Training Order 3 Market-1501 MSMT17 LTCC PRCC SC Average CC Average
AKA [1] 58.3 78.6 3.8 10.1 5.6 12.5 33.3 32.7 31.1 44.4 19.5 22.6
SFT 54.8 77.0 37.0 65.7 16.6 35.5 46.9 48.2 45.9 71.4 31.8 41.9
LwF [33] 63.9 83.0 49.3 75.3 17.8 31.9 46.9 44.6 56.6 79.2 32.4 38.3
CLIP-ReID [12] 63.6 82.7 45.9 73.7 18.8 40.1 49.1 48.7 54.8 78.2 34.0 44.4
Teata (Ours) 77.7 90.7 46.4 73.8 23.2 46.9 58.5 57.3 62.1 82.3 40.9 52.1
Joint-Train 89.1 95.0 70.3 86.6 13.6 28.8 45.8 47.6 79.7 90.8 29.7 38.2
TABLE V: Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 4). The results are reported after the last training phase. The best results are shown in bold.
Training Order 4 LTCC Market-1501 PRCC MSMT17 SC Average CC Average
AKA [1] 11.4 25.0 27.9 51.3 29.6 32.6 11.5 27.1 19.7 39.2 20.5 28.8
SFT 11.5 25.8 53.7 78.1 24.2 23.9 63.5 84.4 58.6 81.3 17.9 24.9
LwF [33] 13.6 25.0 71.3 87.9 31.7 29.1 55.5 79.4 63.4 83.7 22.7 27.1
CLIP-ReID [12] 12.0 24.0 61.1 82.5 23.0 23.3 66.2 86.0 63.7 84.3 17.5 23.7
Teata (Ours) 17.4 37.2 72.4 88.9 48.6 49.6 64.9 83.9 68.7 86.4 33.0 43.4
Joint-Train 13.6 28.8 89.1 95.0 45.8 47.6 70.3 86.6 79.7 90.8 29.7 38.2
TABLE VI: Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 5). The results are reported after the last training phase. The best results are shown in bold.
Training Order 5 LTCC Market-1501 MSMT17 PRCC SC Average CC Average
AKA [1] 8.3 17.9 20.7 42.3 3.3 9.8 33.1 32.5 12.0 26.1 20.7 25.2
SFT 14.4 29.3 61.2 81.4 44.9 72.1 52.0 54.7 53.1 76.8 33.2 42.0
LwF [33] 14.6 28.6 75.8 89.2 48.2 74.3 44.4 41.9 62.0 81.8 29.6 35.3
CLIP-ReID [12] 14.3 33.2 62.3 81.7 41.7 70.0 42.5 42.3 52.0 75.9 28.4 37.8
Teata (Ours) 20.6 41.6 71.4 87.7 53.8 78.6 61.9 62.9 62.6 83.2 41.3 52.3
Joint-Train 13.6 28.8 89.1 95.0 70.3 86.6 45.8 47.6 79.7 90.8 29.7 38.2
TABLE VII: Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 6). The results are reported after the last training phase. The best results are shown in bold.
Training Order 6 LTCC PRCC Market-1501 MSMT17 SC Average CC Average
AKA [1] 11.0 22.4 28.5 30.0 27.6 54.3 10.3 25.5 19.0 39.9 19.8 26.2
SFT 10.6 23.5 22.5 21.1 55.7 79.0 63.6 84.2 59.7 81.6 16.6 22.3
LwF [33] 13.4 25.5 29.7 28.2 71.5 86.6 55.4 79.1 63.5 82.9 16.6 26.9
CLIP-ReID [12] 11.7 23.5 26.2 25.4 61.9 83.2 65.7 85.3 63.8 84.3 19.0 24.5
Teata (Ours) 16.5 36.0 45.4 44.6 75.1 90.0 64.9 84.2 70.0 87.1 31.0 40.3
Joint-Train 13.6 28.8 45.8 47.6 89.1 95.0 70.3 86.6 79.7 90.8 29.7 38.2
IV-C Comparison with the State-of-the-Art
Competitors. We first compare with a widely-used rehearsal-free method AKA [1] for LReID. Then, we reimplement all competitors using CLIP [11] pre-trained ViT-B/16 [53] as the ReID model. Specifically, SFT denotes sequentially fine-tuning the model on each dataset, and Joint-Train denotes training on all datasets jointly, that is, all datasets are available to train the model simultaneously rather than sequentially. Joint-Train is traditionally regarded as the proximate performance upper bound of lifelong learning. We also compare our proposed Teata with LwF [33] and CLIP-ReID [12]. LwF performs knowledge distillation loss on the logit values of old tasks to keep task-specific decision boundaries. CLIP-ReID leverages the CLIP [11] pre-trained text encoder and uses the learned text prompts to regularize the image encoder. CLIP-ReID is designed for conventional person ReID in a stationary domain, and we reimplement it to achieve lifelong learning by retraining text prompts on each sequentially coming dataset.
IV-C1 Results in Seen Domains
[…]
TABLE VIII: Comparison with the state-of-the-art methods in the conventional LReID setting.w/ Ex.” denotes rehearsal-based methods using exemplars. The training order is Market-1501 $\rightarrow$ DukeMTMC-reID $\rightarrow$ CUHK-SYSU $\rightarrow$ MSMT17, and the results are reported after the last training phase. The best results are shown in bold.
@@ IV-C1 Results in Seen Domains
AKA [1] 59.7 80.1 32.7 48.3 82.0 84.4 17.1 34.9 47.9 61.9
GwFReID [35] $\checkmark$ 60.9 81.6 46.7 66.5 81.4 83.9 25.9 52.4 53.7 71.1
LSTKC [38] 56.9 78.0 54.3 70.5 84.4 86.0 34.0 60.3 57.4 73.7
DKP [37] 64.4 83.2 54.4 70.8 85.9 87.6 27.7 51.2 58.1 73.2
KRC [3] $\checkmark$ 60.3 82.3 58.7 72.7 88.9 90.5 43.3 67.7 62.8 78.3
PTKP [2] $\checkmark$ 75.8 89.7 62.0 76.7 85.0 86.3 34.5 60.9 64.3 78.4
LwF [33] 54.0 76.2 64.7 80.7 85.6 87.5 57.7 80.1 65.5 81.1
CLIP-ReID [12] 50.6 74.8 60.0 76.8 84.5 86.4 64.7 84.9 65.0 80.7
@@
Results on LReID-Hybrid benchmarks (Order 1). AKA [1] achieves relatively poor results in the LReID-Hybrid task. Although AKA uses a knowledge graph to represent and accumulate knowledge, operating in the image space makes it difficult to maintain and accumulate knowledge when different tasks exhibit significant differences, especially with clothing state changes. Due to the absence of data from previous domains, SFT suffers greatly from catastrophic forgetting. LwF [33] performs knowledge distillation between current and frozen old models at the cost of sacrificing performance in the current domain. Due to the significant discrepancy across domains caused by variations in clothing states, LwF cannot effectively balance the knowledge from different domains.
Although CLIP-ReID [12] leverages text, it has no obvious advantage compared to LwF, which fine-tunes the pre-trained CLIP image encoder without the supervision of text. The results indicate that the potential of text is still not fully unleashed. In contrast, Teata successfully leverages text semantics to balance and accumulate knowledge from various domains, demonstrating significant superiority over other methods. Note that Teata does not require any task-specific tokens for inference, and the model parameters are shared across all tasks. Strictly speaking, in the setting of LReID-Hybrid, we do not have any clues about the next training task or the testing task, so task-specific tokens are not applicable. Teata outperforms CLIP-ReID [12], which learns text prompts as task-specific tokens at each training task, by a large margin.
@@
It is also worth noting that the performance of Joint-Train is traditionally regarded as the proximate performance upper bound of the lifelong person ReID model. Similarly, when we train the standard and cloth-changing datasets jointly, we find the model can preserve its capability in the same-cloth scenario. However, the results in the cloth-changing domains are far from satisfactory. We argue that the acquired knowledge shows significant discrepancy due to various clothing states across domains. The data imbalance in distribution and quantity between standard and cloth-changing datasets causes the model to be unexpectedly biased toward the former when they are used jointly for training. The results further reveal the challenges faced in our proposed LReID-Hybrid task, which considers clothing variations during the real-world application of person ReID models. Unlike the image model, the text encoder is trained on a massive amount of data, building a compact and well-represented space. Our proposed Teata leverages the text space as an intermediary for knowledge transfer and accumulation, effectively alleviating the representation bias caused by data imbalance.
@@
In Fig. 3, we show the results in each seen domain along the lifelong learning process. Without the help of text semantics, SFT achieves inferior results in each domain. Worse still, when adapting to new domains, it forgets previously acquired knowledge dramatically. Thanks to the consistent and compact text semantics, Teata can accumulate new knowledge while significantly reducing the loss of previously acquired knowledge. We also observe a slight performance improvement on the LTCC dataset when training on PRCC at step 4 compared with step 3. For Teata, the final R-1 accuracy at step 4 is even higher than at step 2. We argue that since both PRCC and LTCC are cloth-changing datasets, the model tends to learn cloth-irrelevant identity representations on them. Therefore, the learned cloth-irrelevant identity clues on PRCC are useful in improving the discriminative ability for pedestrians with clothing changes on LTCC. Teata exhibits a smoother forgetting trend in performance, showcasing its effectiveness in transferring and accumulating knowledge.
[…]
Results on conventional same-cloth LReID benchmarks. Since our proposed method effectively handles the challenging LReID-Hybrid task, it is expected to also perform well when pedestrians do not change their clothes. In Tab. VIII, we compare Teata with state-of-the-art lifelong person ReID methods on the standard LReID benchmarks. Although LSTKC [38] and DKP [37] adopt long short-term knowledge consolidation and distribution-aware prototypes, respectively, they still update the ReID models in the image space, facing the serious knowledge forgetting problem. Exemplars are considered vital in mitigating catastrophic forgetting, and replay-based methods [35, 3, 2] indeed perform better than AKA [1]. However, saving person images can cause privacy issues and is not realistic in real-world applications. It also leads to significant resource consumption with the increase of training domains. We also implement LwF [33] and CLIP-ReID [12] on the conventional LReID benchmarks. Leveraging the superior capability of CLIP, these methods achieve better results than state-of-the-art replay-based methods [35, 3, 2] on the current domain. However, their performance in previous domains is still unsatisfactory. In contrast, Teata achieves a good balance across domains. With the help of consistent text semantics, we beat all competitors by a large margin without relying on impractical or inflexible exemplars.
It is worth noting that when comparing Order 2 of LReID-Hybrid with the conventional same-cloth LReID setting, we find little impact of cloth-changing datasets on same-cloth performance for Teata. This demonstrates its effectiveness in handling knowledge mismatch and alleviating the knowledge forgetting caused by varying clothing states.
@@
TABLE IX: Unseen domain comparisons with other competitors trained in the LReID-Hybrid setting (Order 1). The results are given by the trained model after the last training step. $\dagger$” means the light version of Celeb-reID. The best results are shown in bold.
@@
TABLE X: Unseen domain comparisons with state-of-the-art models trained in the conventional LReID setting. The training order is Market-1501 $\rightarrow$ DukeMTMC-reID $\rightarrow$ CUHK-SYSU $\rightarrow$ MSMT17. The results are given by the trained model after the last training step. $\dagger$” indicates the light version of Celeb-reID. The best results are shown in bold.
@@
PTKP [2] 71.1 69.4 66.4 66.1 33.9 25.3 53.9 46.1 38.4 27.3 24.2 21.8 5.3 9.2 52.7 46.8 14.8 15.5
KRC [3] 79.1 78.0 70.0 69.0 33.6 25.5 51.3 44.0 51.8 40.6 20.2 17.5 6.3 12.2 57.2 51.4 13.3 14.9
LwF [33] 76.5 74.7 67.2 66.7 36.8 28.0 56.2 47.0 58.0 46.4 35.0 33.5 6.6 13.2 58.9 52.6 20.8 23.4
CLIP-ReID [12] 77.2 75.6 65.3 64.9 35.4 25.3 56.7 48.5 53.3 40.5 36.4 34.9 6.0 11.4 57.6 51.0 21.2 23.2
@@
IV-C2 Results in Unseen Domains